OpenStudy (anonymous):

Make: r the subject of V = 4/3 (fraction) * pi * r^3

OpenStudy (anonymous):

3V = 4 pi r^3 (3V)/(4pi) = r^3 r^(1/3) = r =( (3V)/(4pi) )^(1/3)